# leftmargin=0pt, rightmargin=0pt, backgroundcolor=bggray, middlelinecolor=black, roundcorner=3 *[subfigure]position=bottom

### v1

@@ Abstract.
AI-based systems such as language models can replicate and amplify social biases reflected in their training data. Among other questionable behavior, this can lead to LM-generated text–and text suggestions–that contain normatively inappropriate stereotypical associations. In this paper, we consider the question of how “debiasing” a language model impacts stories that people write using that language model in a predictive text scenario. We find that ($n=414$), in certain scenarios, language model suggestions that align with common social stereotypes are more likely to be accepted by human authors. Conversely, although anti-stereotypical language model suggestions sometimes lead to an increased rate of anti-stereotypical stories, this influence is far from sufficient to lead to “fully debiased” stories.
@@
ccs: Human-centered computing Empirical studies in HCIccs: Computing methodologies Machine learning
Figure 1. We study human-AI co-writing with biased predictive text models. In settings like the one pictured, we find that though anti-stereotypical suggestions significantly decrease the amount of pro-stereotypical stories written, this is not enough to remove (let alone reverse) the pro-stereotypical bias in the co-written stories.
[…]
Examples of people writing with no suggestions, pro-stereotypical suggestions and anti-stereotypical suggestions. All start from the text “In the first 100 days, the new president was determined to focus”. In the no suggestions and pro-stereotypical cases, the person continues with the word “his”, but with anti-stereotypical suggestions the person notes that the suggested “her” also works here, even if it wasn’t what they were originally thinking. On the right, we show the gender breakdown of president characters written by our study participants with these kinds of suggestions. We see that despite participants writing masc-coded presidents significantly less often and fem-coded presidents marginally more often when shown anti-stereotypical suggestions over no suggestions, they still write fem-coded presidents significantly less often than masc-coded
@@ 1. Introduction
Predictive text systems have become a commonly used tool in human communication, with $44\%$ of Americans reporting using predictive text at least somewhat often. [^1] https://civicscience.com/ai-in-daily-life-people-increasingly-embrace-predictive-text/ While users and developers may see predictive text technology as producing “neutral” output, it is well known that the language models that underlie predictions often pick up on—and even amplify—social biases, including those present in their training data (Gupta et al., 2024) as well as those due to structural factors around their creation (Blodgett et al., 2020). These language model biases can directly lead to the generation of text that causes representational harms to users (Barocas et al., 2017) including alienation, erasure, and—the topic of this paper—stereotyping.
@@
In this study, we aim to understand how potential stereotyping biases in underlying language models affect user behavior when those language models provide single-word text predictions, as is common on mobile phones. In a pre-registered  [^2] https://aspredicted.org/SHD_PM4 and IRB approved [^3] IRB identifier information removed for anonymization online study ($n=414$), we asked participants to write short English stories with (treatment condition) or without (control condition) the help of a predictive text system. In the treatment condition, when the participants were provided with text predictions, these predictions were generated by either a language model that was designed to make suggestions that aligned with social stereotypes, or on one that was designed to challenge social stereotypes surrounding gender and sexuality. These stereotypes included gender-occupation stereotypes (pro-stereotypical: a doctor who uses he/him pronouns; anti-stereotypical: a doctor who uses she/her pronouns) as well as personality stereotypes based on the Agency-Belief-Communion (ABC) model from social psychology (Koch et al., 2016) (pro-stereotypical: men are untrustworthy; anti-stereotypical: women are untrustworthy).
@@
Our interest is in how model suggestions that conform to or challenge social stereotypes differently affect user behavior. We measure—in the stories written with or without model suggestions—how often the overall stories that people write conform to (vs challenge) stereotypes, as well as how often participants make use of the text predictions in the pro-stereotypical vs anti-stereotypical conditions. While much work has been done to reduce stereotypes and biases in language models themselves (Gupta et al., 2024), we are not aware of prior work that investigates how this impacts the writing of people who use those systems. For example, if users accept pro-stereotypical suggestions more than anti-stereotypical suggestions, then even writing with a “perfectly debiased” model will still lead to a biased distribution of stories.
@@
Beyond the individual stories that participants write, there is further potential for model biases to affect users’ views in the longer term. A standard model in social psychology connects stereotypes—over-generalized views about a group—directly to the formation of prejudices—the believes one holds about a group—and from there to discrimination—actions against a group (Fiske, 1998). Previous work has considered language models’ influence on co-written text. For example, Arnold et al. (2018) and Bhat et al. (2021, 2023) found that co-writing with a biased language model can affect users’ expressed sentiment in reviews. Jakesch et al. (2023), Dhillon et al. (2024), and Padmakumar and He (2024) found that co-writing can affect the position and diversity of the views users express on topics including the societal impact of social media, whether college athletes should be paid, etc.
In our study, we find that in certain scenarios, participants are more likely to accept predictive text suggestions that align with gender stereotypes. When considering the gender distribution of co-written characters, we see that anti-stereotypical suggestions (such as suggesting a character in a male-dominated profession is a woman) lead people to write marginally more anti-stereotypical stories than they do without predictive text suggestions. However, this increase is never enough to reverse the trend toward pro-stereotypical stories. For example, although a predictive text system that suggests that a president character is a woman marginally increases the proportion of fem-coded presidents in stories, participants nonetheless write far more stories with masc-coded presidents than fem-coded presidents.
When considering the ABC traits assigned to characters of different genders, our findings are mixed. For example, we find evidence that predictive text suggestions that a masc-coded teacher character is “likeable” marginally increases the proportion of “likeable” masc-coded teachers in stories. However, participants preferred to write “likeable” teachers regardless of gender. Overall, we still note many writing scenarios where participants are significantly more likely to accept pro-stereotypical predictive text suggestions than anti-stereotypical.
@@ 2. Related Work
Humans are not free from biases and stereotypes (Greenwald and Banaji, 1995). People have been found to evaluate identical work in academic settings more favorably when attributed to male authors compared to female authors (Moss-Racusin et al., 2012; Goldberg, 1968). And implicit gender biases in promotion committees have been linked to lower advancement rates for women in STEM fields, especially when committees fail to recognize external barriers faced by women (Régner et al., 2019).
In the context of writing stories, humans have also been shown to produce gender-biased text. Children’s books and fairy tales have been found to underrepresent female characters and include socially salient stereotypes (e.g., Toro Isaza et al., 2023; Filipović, 2018; Lewis et al., 2022; Gooden and Gooden, 2001). Toro Isaza et al. (2023) analyze gender differences in the kinds of events fairy tale characters participate in throughout a narrative arc. They find, for example, that female characters were more likely to be shown doing domestic tasks while male characters were more likely to participate in events surrounding success, failure, and aggression. Prior work has also found that the gender stereotypes present or absent in “the reading materials to which we expose children shape their attitudes, their understanding and their behavior” affecting their “self-concept, potential for achievement and perceptions of others” and stereotypical beliefs and attitudes (Peterson and Lach, 1990).
Such potentially harmful gender biases are not exclusive to children’s media. For example, men are more represented than women in commercial films in terms of time speaking (Google, 2017) and time spent on-screen (Jang et al., 2019). Books written by men spend about a third to a fourth of the space describing characters on describing female characters, while books written by women are closer to equal (Underwood et al., 2018). While how gender is represented in literature has changed over time, gendered differences in how characters are described (especially physically) are present even in more modern literature (Underwood et al., 2018). Tropes in media also reveal gender bias, with highly male-associated tropes covering topics such as “money and strength” and highly female-associated tropes covering topics such as “motherhood and pregnancy” (Gala et al., 2020).
Our work concerns how gender biases like can potentially be exacerbated by writing with a biased predictive text system. Exposure to these biases, both for the authors using the predictions and the readers consuming the final result, may affect stereotypical beliefs and perceptions, especially for younger, more impressionable audiences.
@@
Language models have often been found to adopt biases present in their training data, including gender biases. Much of the work on gender bias in models focuses on intrinsic biases (e.g., Bolukbasi et al., 2016; Caliskan et al., 2017)—biases present in internal model representations such as word embedding vectors—or extrinsic biases—biases in downstream task performance such as summarization or question answering (e.g., Parrish et al., 2022).
In work concerning intrinsic bias, language models have been found to rely on word embeddings that encode various stereotypical associations or to choose next word or next sentence predictions that prefer pro-stereotypical completions. Such studies have demonstrated intrinsic biases covering associations between gender and occupation (Bolukbasi et al., 2016; Zhao et al., 2019), gender and arts vs science/math (Caliskan et al., 2017; Guo and Caliskan, 2021; Kurita et al., 2019), and gender and traits like “polite” or “burly” (Nadeem et al., 2021), “trustworthy” vs “untrustworthy” (Cao et al., 2022b). Other work has found evidence of intrinsic anti-queer biases in models such as assigning sentences about queer couples a lower pseudo-log-likelihood than minimally edited sentences about queer couples (Nangia et al., 2020) or sentences containing stereotypes about the queer community a higher pseudo-log-likelihood than minimally edited sentences about straight people (Felkner et al., 2023).
In work concerning extrinsic bias, models have been found to over-rely on gender stereotypes, gendered associations, etc on various downstream tasks such as coreference resolution (Rudinger et al., 2018; Zhao et al., 2018; Webster et al., 2018; Cao and Daumé, 2021; Baumler and Rudinger, 2022), sentiment analysis (Kiritchenko and Mohammad, 2018), emotion attribution (Plaza-del Arco et al., 2024), occupation classification (De-Arteaga et al., 2019), question answering (Parrish et al., 2022), etc, leading to poorer performance on examples that do not match gender stereotypes. For example, models over-rely on gender-occupation stereotypes in coreference resolution, even in light of syntactic structures or common-sense information which should make the correct answer clear (Rudinger et al., 2018; Zhao et al., 2018). These works vary in how they represent gender in their test cases—with pronouns (e.g., Rudinger et al., 2018), gender-associated names (e.g., An et al., 2023), gender-associated terms like “woman” or “daughter” (e.g., Bolukbasi et al., 2016), etc.
These intrinsic and extrinsic measures do not always correlate (Cao et al., 2022a; Goldfarb-Tarrant et al., 2021), meaning that just because a bias is present or absent for a given intrinsic measure, this does not mean the users will or will not experience biased outcomes when using the model for a downstream task.
@@
Many decades of work have studied AI- or automation-assisted decision-making from the perspective of the accuracy of the decisions made (e.g., Mosier et al., 1996; Parasuraman and Riley, 1997; Lee and See, 2004). Here, we are interested in how the bias of a human-AI assemblage relates to the bias of humans-alone or AI-alone. Prior work on human-AI decision-making has found that the bias of a human-AI team is not simply equal to the sum of its parts and can depend on factors such as the decision-making task and whether or how the AI’s suggestions are justified (e.g., Schoeffer et al., 2024; Goyal et al., 2024; Wang et al., 2023; Peng et al., 2022; De-Arteaga et al., 2020). Our paper considers the task of human text authorship with the help of word-level suggestions given by a predictive text system. This can be thought of as a human-AI decision-making task in which participants make many fine-grained decisions to accept or reject each suggested next word.
De-Arteaga et al. (2020) study how model suggestions affect decisions to screen in child welfare services calls for further investigation. They find that model recommendations not only increase the overall quality of decisions, but they decrease the gap in screen-in rates for White and Black children showing there was not a “difference in willingness to adhere to the recommendation that would compound previous racial injustices.”
@@
Wang et al. (2023) assess how making decisions with a biased AI affects the fairness of decisions in how much to bid on a rental house. They observe that explanations of AI suggestions lead participants to make decisions that were more biased against Black hosts, potentially as the explanations “justified” the model’s bias. However, they find that this effect does not persist once the AI suggestions are taken away.
@@
In our paper, we examine the effect of social biases in predictive text suggestions on co-writing. Writing with predictive text is a task that many laypeople encounter in their day-to-day lives. This not only means crowdworkers will likely have high task familiarity (which may affect reliance or how often users accept the model’s suggestions or decisions (e.g., Wang and Yin, 2022)) but also that the influences identified in the study are applicable to a large portion of the population. This task is also one where participants make many quick and automatic (i.e., System 1 (Kahneman, 2011)) decisions, making it a good surrogate task for stereotypes and implicit biases.
@@
Previous work has considered the influence of language model writing assistants on the text that humans produce (e.g., Jakesch et al., 2023; Bhat et al., 2021; Arnold et al., 2020; Padmakumar and He, 2024; Mirowski et al., 2023; Bhat et al., 2023; Arnold et al., 2018; Buschek et al., 2021; Dhillon et al., 2024).
Arnold et al. (2018) and Bhat et al. (2021, 2023) consider how predictive text can bias the sentiment of users’ writing. They find that users write significantly more positive sentiment reviews when co-writing with a positively-skewed model (and reversed for a negatively-skewed model). Jakesch et al. (2023) find similar results in the context of argumentative essay writing. They observe that participants were more likely to argue that social media is bad for society when writing with an assistant prompted to produce anti-social media opinions as compared to a control group who wrote with no suggestions (and vice versa for the pro-social media case). Dhillon et al. (2024) similarly find that AI suggestions in co-writing can influence users’ opinions, especially when the AI provides longer, paragraph-level suggestions. Padmakumar and He (2024) also consider the context of argumentative writing, finding that writing with different language model assistants leads to measurably different levels of homogeneity in essays, depending on how diverse the suggestions are from the underlying models.
While these studies all consider the influence of model suggestions on writing, the form of these suggestions is variable—ranging from a single word (e.g., Arnold et al., 2020) to an entire paragraph (e.g., Jakesch et al., 2023). Our work specifically examines the impact of word-level suggestions. Prior research has found that longer suggestions may increase impact of AI suggestions users’ expressed opinions (Dhillon et al., 2024). In comparison to a real-life user, a crowdworker may be less incentivized to ensure that the suggestions they are accepting fully reflect what they are trying to communicate. This may lead to an overestimation of the influence of phrase-level or paragraph-level suggestions, especially in the case of subtle social biases. For example, Macrae et al. (1994) found that stereotypes serve as “cognitive shortcuts” that facilitate quicker decision-making at the cost of decreased accuracy and lower levels of fairness.
@@
3. Hypotheses

### v2

# by*[subfigure]position=bottom

@@ Abstract.
AI-based systems such as language models have been shown to replicate and even amplify social biases reflected in their training data. Among other questionable behaviors, this can lead to AI-generated text–and text suggestions–that contain normatively inappropriate stereotypical associations. Little is known, however, about how this behavior impacts the writing produced by people using these systems. We address this gap by measuring how much impact stereotypes or anti-stereotypes in English single-word LM predictive text suggestions have on the stories that people write using those tools in a co-writing scenario. We find that ($n=414$), LM suggestions that challenge stereotypes sometimes lead to a significantly increased rate of anti-stereotypical co-written stories. However, despite this increased rate of anti-stereotypical stories, pro-stereotypical narratives still dominated the co-written stories, demonstrating that technical debiasing is only a partially effective strategy to alleviate harms from human-AI collaboration.
@@
journalyear: 2026copyright: ccconference: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems; April 13–17, 2026; Barcelona, Spainbooktitle: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems (CHI ’26), April 13–17, 2026, Barcelona, Spaindoi: 10.1145/3772318.3790733isbn: 979-8-4007-2278-3/2026/04ccs: Human-centered computing Empirical studies in HCIccs: Computing methodologies Machine learning
Figure 1. We study human-AI co-writing with biased predictive text models. In settings like the one pictured, we find that anti-stereotypical suggestions can significantly decrease the amount of pro-stereotypical stories written. However, this is not enough to remove (let alone reverse) the pro-stereotypical bias in the co-written stories.
Left: three example writing interactions under no-suggestion, pro-stereotypical, and anti-stereotypical conditions, all starting from “In the first 100 days, the new president was determined to focus.” In the first two conditions the participant continues with “his,” while in the anti-stereotypical condition the participant considers the suggested “her.” Right: bar chart showing the gender distribution of president characters written by participants across the three conditions. The chart includes a highlighted paragraph and arrows stating that participants write masculine-coded presidents “significantly less often” and feminine-coded presidents “marginally more often” with anti-stereotypical suggestions, but still write feminine-coded presidents “significantly less often” than masculine-coded.
@@ 1. Introduction
Predictive text systems have become a commonly used tool in human communication, with $44\%$ of Americans reporting using predictive text at least somewhat often. [^1] https://civicscience.com/ai-in-daily-life-people-increasingly-embrace-predictive-text/ While users and developers may see predictive text technology as producing “neutral” output, it is well known that the language models that underlie predictions often pick up on—and even amplify—social biases, including those present in their training data (Gupta et al., 2024) as well as those due to structural factors around their creation (blodgett-etal-2020-language). These language model biases can directly lead to the generation of text that causes representational harms to users (Barocas et al., 2017; dev-etal-2022-measures) including alienation, erasure, disparagement, dehumanization and—the topic of this paper—stereotyping. Our work focuses on distribution-based stereotyping (Katzman et al., 2023)—in the stories written with or without model suggestions—how often the overall stories that people write conform to (vs challenge) stereotypes, as well as how often participants make use of the text predictions in the pro-stereotypical vs anti-stereotypical conditions.
@@
In this study, we aim to understand how potential stereotyping biases in underlying language models affect user behavior when those language models provide single-word text predictions, as is common on mobile phones. In a pre-registered  [^2] https://aspredicted.org/SHD_PM4 and IRB approved [^3] University of Maryland IRB package number 2099750-2 online study ($n=414$), we asked participants to write short English stories with (treatment condition) or without (control condition) the help of a predictive text system. In the treatment condition, when the participants were provided with text predictions, these predictions were generated—on a per-story basis—by either a language model that was designed to make suggestions that aligned with social stereotypes, or one that was designed to challenge social stereotypes surrounding gender and sexuality. These stereotypes included gender-occupation stereotypes (pro-stereotypical: a doctor who uses he/him pronouns; anti-stereotypical: a doctor who uses she/her pronouns) as well as personality stereotypes based on the Agency-Belief-Communion (ABC) model from social psychology (Koch et al., 2016) (pro-stereotypical: men are untrustworthy; anti-stereotypical: women are untrustworthy). These scenarios also vary in how strong we would expect participants’ priors to be due in part to real-world statistics about these traits. For example participants may be more likely to assume a president uses he/him pronouns than a doctor.
@@
Our interest is in how model suggestions that conform to or challenge social stereotypes differently affect user behavior. While much work has been done to reduce stereotypes and biases in language models themselves (Gupta et al., 2024), we are not aware of prior work that investigates how this debiasing impacts the writing of people who use those systems. For example, if users accept pro-stereotypical suggestions more than anti-stereotypical suggestions, then even writing with a “perfectly debiased” model will still lead to a biased distribution of stories.
@@
Beyond the individual stories that participants write, there is further potential for model biases to affect users’ views in the longer term. A standard model in social psychology connects stereotypes—over-generalized views about a group—directly to the formation of prejudices—the beliefs one holds about a group—and from there to discrimination—actions against a group (Fiske, 1998). Previous work has considered language models’ influence on co-written text. For example, Arnold et al. (2018) and bhat-etal-2021-people; Bhat et al. (2023) found that co-writing with a biased language model can affect users’ expressed sentiment in reviews while Agarwal et al. (2025) found that co-writing with a Western-centric model can encourage adoption of Western writing and cultural norms. Jakesch et al. (2023), Dhillon et al. (2024), and Padmakumar and He (2024) found that co-writing can affect the position and diversity of the views users express on topics including the societal impact of social media, whether college athletes should be paid, etc.
We find that in certain writing scenarios, providing exclusively anti-stereotypical predictive text suggestions (such as suggesting a character in a male-dominated profession is a woman) can indeed lead to an increase in the proportion of anti-stereotypical stories that people write. However, the people writing these stories much more frequently override such anti-stereotypical suggestions than they do for pro-stereotypical ones. In fact, even when all AI suggestions were anti-stereotypical, we do not observe any scenarios where participants wrote more anti-stereotypical stories than pro-stereotypical (as illustrated in the “president” writing scenario, Figure 1). Since a system that provides exclusively anti-stereotypical suggestions is unlikely to be deployed in practice, our results should be viewed as an upper bound on how much purely technical “debiasing” can potentially mitigate stereotypical writing in human-AI co-writing scenarios. The effect of a more “realistically” debiased predictive text model (e.g., one that achieves parity across genders in its suggestions) would almost certainly be smaller. For developers and practitioners who wish to encourage a more “fair” distribution of stereotype-relevant content in human-AI written stories, we conclude that while “debiasing” the model may have some positive impact on human behavior, such suggestions alone may be an insufficient intervention.
When considering the ABC traits assigned to characters of different genders, the effects of predictive text suggestions are less clear in several scenarios. In some scenarios, this may be because the ABC traits are less often or less clearly marked in treatment ($54\%\pm 16$ on average across scenarios) and control writing than gender ($70\%\pm 20$ on average across scenarios), and some of the studied gender-trait associations may be weaker than gender-occupation associations. Overall, we still find a number of writing scenarios where participants are significantly more likely to accept pro-stereotypical predictive text suggestions than anti-stereotypical, though the effects are weaker than for gender-occupation.
@@ 2. Related Work
Humans are not free from biases and stereotypes (Greenwald and Banaji, 1995). People have been found to evaluate identical work in academic settings more favorably when attributed to male authors compared to female authors (Moss-Racusin et al., 2012; Goldberg, 1968). And implicit gender biases in promotion committees have been linked to lower advancement rates for women in STEM fields, especially when committees fail to recognize external barriers faced by women (Régner2019).
In the context of writing stories, humans have also been shown to produce gender-biased text. Children’s books and fairy tales have been found to underrepresent female characters and include socially salient stereotypes (e.g., toro-isaza-etal-2023-fairy; Filipović, 2018; Lewis et al., 2022; Gooden and Gooden, 2001). toro-isaza-etal-2023-fairy analyze gender differences in the kinds of events fairy tale characters participate in throughout a narrative arc. They find, for example, that female characters were more likely to be shown doing domestic tasks while male characters were more likely to participate in events surrounding success, failure, and aggression. Prior work has also found that the gender stereotypes present or absent in “the reading materials to which we expose children shape their attitudes, their understanding and their behavior” affecting their “self-concept, potential for achievement and perceptions of others” and stereotypical beliefs and attitudes (Peterson and Lach, 1990).
Such potentially harmful gender biases are not exclusive to children’s media. For example, men are more represented than women in commercial films in terms of time speaking (Google, 2017) and time spent on-screen (Jang et al., 2019). Books written by men spend about a third to a fourth of the space describing characters on describing female characters, while books written by women are closer to equal (Underwood et al., 2018). While how gender is represented in literature has changed over time, gendered differences in how characters are described (especially physically) are present even in more modern literature (Underwood et al., 2018). Tropes in media also reveal gender bias, with highly male-associated tropes covering topics such as “money and strength” and highly female-associated tropes covering topics such as “motherhood and pregnancy” (gala-etal-2020-analyzing).
Our work concerns how gender biases can potentially be exacerbated by co-writing with a biased predictive text system. Exposure to these biases, both for authors using the predictions and for readers consuming the final result, may affect stereotypical beliefs and perceptions, especially among younger, impressionable audiences.
@@
Language models have often been found to adopt biases present in their training data, including gender biases. Much of the work on gender bias in models focuses on intrinsic biases (e.g., Bolukbasi et al., 2016; Caliskan et al., 2017)—biases present in internal model representations such as word embedding vectors—or extrinsic biases—biases in downstream task performance such as summarization or question answering (e.g., parrish-etal-2022-bbq).
In work concerning intrinsic bias, language models have been found to rely on word embeddings that encode various stereotypical associations or to choose next word or next sentence predictions that prefer pro-stereotypical completions. Such studies have demonstrated intrinsic biases covering associations between gender and occupation (Bolukbasi et al., 2016; zhao-etal-2019-gender; An et al., 2025), gender and arts vs science/math (Caliskan et al., 2017; Guo and Caliskan, 2021; kurita-etal-2019-measuring), and gender and traits like “polite” or “burly” (nadeem-etal-2021-stereoset), “trustworthy” vs “untrustworthy” (cao-etal-2022-theory). Other work has found evidence of intrinsic anti-queer biases in models such as assigning sentences about queer couples a lower pseudo-log-likelihood than minimally edited sentences about queer couples (nangia-etal-2020-crows) or sentences containing stereotypes about the queer community a higher pseudo-log-likelihood than minimally edited sentences about straight people (felkner-etal-2023-winoqueer).
In work concerning extrinsic bias, models have been found to over-rely on gender stereotypes and gendered associations on downstream tasks such as coreference resolution (rudinger-etal-2018-gender; zhao-etal-2018-gender; webster-etal-2018-mind; Cao and Daumé, 2021; baumler-rudinger-2022-recognition), sentiment analysis (kiritchenko-mohammad-2018-examining), emotion attribution (Plaza-del-Arco et al., 2024), occupation classification (De-Arteaga et al., 2019), question answering (parrish-etal-2022-bbq), leading to poorer performance on examples that do not match gender stereotypes. For example, models over-rely on gender-occupation stereotypes in coreference resolution, even in light of syntactic structures or common-sense information which should make the correct answer clear (rudinger-etal-2018-gender; zhao-etal-2018-gender). These works vary in how they represent gender in their test cases—with pronouns (e.g., rudinger-etal-2018-gender), gender-associated names (e.g., an-etal-2023-sodapop), gender-associated terms like “woman” or “daughter” (e.g., Bolukbasi et al., 2016), etc.
These intrinsic and extrinsic measures do not always correlate (cao-etal-2022-intrinsic; goldfarb-tarrant-etal-2021-intrinsic), meaning that just because a bias is present or absent for a given intrinsic measure, this does not mean the users will or will not experience biased outcomes when using the model for a downstream task.
@@
Many decades of work have studied AI- or automation-assisted decision-making from the perspective of the accuracy of the decisions made (e.g., Mosier et al., 1996; Parasuraman and Riley, 1997; Lee and See, 2004). Here, we are interested in how the bias of a human-AI assemblage relates to the bias of humans-alone or AI-alone. Prior work on human-AI decision-making has found that the bias of a human-AI team is not simply equal to the sum of its parts and can depend on factors such as the decision-making task and whether or how the AI’s suggestions are justified (e.g., Schoeffer et al., 2024; Goyal et al., 2024; Wang et al., 2023b; Peng et al., 2022; De-Arteaga et al., 2020; Zipperling et al., 2025). Our paper considers the task of human text authorship with the help of word-level suggestions given by a predictive text system. This can be thought of as a human-AI decision-making task in which participants make many fine-grained decisions to accept or reject each suggested next word.
De-Arteaga et al. (2020) study how model suggestions affect decisions to screen in child welfare services calls for further investigation. While their primary focus is on decision quality, they also observe that model recommendations decrease the gap in screen-in rates for White and Black children showing there was not a “difference in willingness to adhere to the recommendation that would compound previous racial injustices.”
@@
Wang et al. (2023b) assess how making decisions with a biased AI affects the fairness of decisions in how much to bid on a rental house. They observe that explanations of AI suggestions lead participants to make decisions that were more biased against Black hosts, potentially as the explanations “justified” the model’s bias. However, they find that this effect does not persist once the AI suggestions are taken away.
@@
While these previous works focus on how biases in AI suggestions affect the decisions of a human-AI team, other studies have focused on the effects of collaborating with a “debiased” system. Krause et al. (2025) and Wang et al. (2023a, 2022) consider the effect of debiased AI suggestions in the context of college major and career recommendations. They find that participants overall prefer gender biased suggestions with Krause et al. (2025) noting a stronger effect in female participants. Zipperling et al. (2025) consider the effect of “alignment” between human and AI bias more generally. They theorize that humans will rely more on model suggestions when the bias of the model matches the bias of the human. They find that participants who produce more gender-biased decisions alone are more likely to rely on a “gendered” AI than an “ungendered” AI.
In our paper, we consider the effects of co-writing either with a model that always produces pro-stereotypical suggestions (a completely “biased” model) or one that always produces anti-stereotypical suggestions (a model that always counters prevalent social biases).
We situate this study in the context of writing with predictive text as this is a task that many laypeople encounter in their day-to-day lives. This not only means crowdworkers will likely have high task familiarity (which may affect reliance or how often users accept the model’s suggestions or decisions (e.g., Wang and Yin, 2022)) but also that the influences identified in the study are applicable to a large portion of the population. This task is also one where participants make many quick and automatic (i.e., System 1 (Kahneman, 2011)) decisions, making it a good surrogate task for stereotypes and implicit biases.
@@
Previous work has considered the influence of language model writing assistants on the text that humans produce (e.g., Jakesch et al., 2023; bhat-etal-2021-people; Arnold et al., 2020; Padmakumar and He, 2024; Mirowski et al., 2023; Bhat et al., 2023; Arnold et al., 2018; Buschek et al., 2021; Dhillon et al., 2024; Agarwal et al., 2025).
Arnold et al. (2018) and bhat-etal-2021-people; Bhat et al. (2023) consider how predictive text can bias the sentiment of users’ writing. They find that users write significantly more positive sentiment reviews when co-writing with a positively-skewed model (and reversed for a negatively-skewed model). Jakesch et al. (2023) find similar results in the context of argumentative essay writing. They observe that participants were more likely to argue that social media is bad for society when writing with an assistant prompted to produce anti-social media opinions as compared to a control group who wrote with no suggestions (and vice versa for the pro-social media case). Dhillon et al. (2024) similarly find that AI suggestions in co-writing can influence users’ opinions, especially when the AI provides longer, paragraph-level suggestions. Padmakumar and He (2024) also consider the context of argumentative writing, finding that writing with different language model assistants leads to measurably different levels of homogeneity in essays, depending on how diverse the suggestions are from the underlying models. Agarwal et al. (2025) further find that co-writing with an AI system can homogenize writing towards particularly toward Western cultural norms leading, for example, Indian authors use more generic or exoticized descriptions of Indian festivals and foods.
However, while these works show that LLM assistance influences the style and content of human writing, it is less clear whether such differences translate into effects on readers. For instance, Biswas et al. (2025) find that while prior experience using an LLM in a low-resource language affects their reliance on LLM suggestions when co-writing in English, these differences do not affect the downstream persuasiveness of co-written text.
While these studies all consider the influence of model suggestions on writing, they differ in the form of these suggestions—ranging from a single word (e.g., Arnold et al., 2020) to an entire paragraph (e.g., Jakesch et al., 2023). Our work specifically examines the impact of word-level suggestions. Prior research has found that longer suggestions may increase impact of AI suggestions users’ expressed opinions (Dhillon et al., 2024). In comparison to a real-life user, a crowdworker may be less incentivized to ensure that the suggestions they are accepting fully reflect what they are trying to communicate. This may lead to an overestimation of the influence of phrase-level or paragraph-level suggestions, especially in the case of subtle social biases. For example, Macrae et al. (1994) found that stereotypes serve as “cognitive shortcuts” that facilitate quicker decision-making at the cost of decreased accuracy and lower levels of fairness.
@@
3. Research Question and Hypotheses
Prior work has shown that stereotypes in humans can be deeply held and resistant to change, and that AI models can encode similar human-like biases and stereotypes (See section 2). While existing literature demonstrates that AI suggestions can influence aspects of co-writing, such as sentiment and opinions, it remains unexplored whether and how stereotypes—often harder to meaningfully change than other, more malleable opinions—might specifically impact co-writing through predictive text. Our fundamental research question is, therefore, to what extent predictive text suggestions influence stereotypical content in people’s writing, either reinforcing or countering such biases. Although AI suggestions may influence writing in certain ways, they may not effectively nudge writing away from deeply rooted human biasses.